OpenStudy (anonymous):

solve using quad formula 2x^2+12x=-7

OpenStudy (anonymous):

We have the equation: \[2x^2+12x=-7\] How the quadratic formula works is: "Give me a equation in the form of \(ax^2+bc+c=0\) and I will tell you that the roots, (or the solutions) of the equation are exactly: \(x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)" So we need to re-arrange the equation \(2x^2+12x=-7\) into that form. It's really not that bad as adding 7 to both sides will result in: \(2x^2+12x+7=0\). So now that we have it in that form, lets identify our variables. We have the variables x, a, b, and c. We have a formula which will give us x if we give it a, b, and c. So lets identify those three variables. We can tell just by looking at it that: \[\eqalign{ &a=2\\ &b=12\\ &c=7\\ }\] So put those into the formula: \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] And tell me what you get! :)

OpenStudy (anonymous):

i did and i got x= -12√200 / 4 i couldnt type it as good as you did but u get what im saying anyways thats were im stuck

OpenStudy (anonymous):

Well lets see: \[\eqalign{ x&=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \\ &=\frac{-12\pm\sqrt{144-4(2)(7)}}{4} \\ &=\frac{-12\pm\sqrt{144-64}}{4} \\ &=\frac{-12\pm\sqrt{80}}{4} \\ &=\frac{-12\pm\sqrt{16\times5}}{4} \\ &=\frac{-12\pm\sqrt{16}\sqrt{5}}{4} \\ &=\frac{-12\pm4\sqrt{5}}{4} \\ &=-3\pm\sqrt{5} \\ }\] So that means that our answers are either: \[x=-3+\sqrt{5}\approx0.76\phantom{......}or\phantom{......}x=-3-\sqrt{5}\approx-5.23\]

OpenStudy (anonymous):

Err correction...the first one is \(-0.76\), not \(0.76\)

OpenStudy (anonymous):

Thanks a lot! i cant believe how simple of a mistake i was making and i couldnt notice it! that what happens when you study and are sleepy lol

OpenStudy (anonymous):

Haha I know eh? Anyways, no problem!